OpenStudy (anonymous):

I've been stuck on this problem for a long time. How do you do this problem? What's the answer? Solve for x: 3/4 x + 5/8 = 4x x =_____ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar. (2 points)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\frac{3}{4}x+\frac{5}{8}=4x\]?

OpenStudy (misty1212):

easiest thing to do is multiply both sides by 8 to get rid of annoying fractions and start with \[6x+5=16x\] solve that one in two steps

OpenStudy (anonymous):

what's the final answer though?

OpenStudy (misty1212):

idk i didn't do it i am sure you can solve \[6x+5=16x\]

OpenStudy (anonymous):

is the answer x= 1/2